# Adaptive Data-Knowledge Alignment in Genetic Perturbation Prediction

### v1

@@ 1 Introduction
Two complementary approaches have emerged, either by leveraging latent representations trained on extensive cell data (Lotfollahi et al., 2023; Theodoris et al., 2023; Cui et al., 2024; Hao et al., 2024) or incorporating prior biological knowledge for inductive biases (Roohani et al., 2024; Wang et al., 2024; Littman et al., 2025; Wenkel et al., 2025). Yet, both approaches provide limited insights into the biological mechanisms underlying their predictions. Data-driven models operate as black boxes, making it difficult to understand which regulatory relationships drive specific predictions (Bendidi et al., 2024). While hybrid methods incorporate prior biological knowledge, they treat this knowledge as static constraints rather than interpretable and updatable representations of biological understanding. Importantly, current approaches provide no end-to-end solution to identify and resolve divergences between data-driven learning and existing knowledge, which limits opportunities for continual refinement of biological understanding. (Gavriilidis et al., 2024; Kedzierska et al., 2025).
@@
  • Adaptive Neuro-Symbolic Alignment. We align neural and symbolic predictions from inconsistent information sources by adaptively weighting neural and symbolic components with a gradient-free optimization mechanism.
@@
  • Results. ALIGNED outperforms existing methods in balanced consistency with both data and knowledge. In addition, ALIGNED’s knowledge refinement can re-discover cross-referenced regulatory relationships. Our results demonstrate effective translation from prediction to enhanced mechanistic interpretation.
@@ 2.1 Problem Setting
We formalize the prediction of genome-scale response to genetic perturbation as a ternary classification problem. The goal is to learn a function $f:\{-1,0,1\}^{n}\rightarrow\{-1,0,1\}^{n}$, where $n$ is the total number of genes. The input values 1-1, 0, and $1$ represent negative perturbation (deletion or knockout), no perturbation, and positive perturbation (overexpression), respectively. The output values indicate decreased expression, no significant change, or increased expression for each gene.
@@ 2.2 Symbolic reasoning over gene regulatory networks
We explain symbolic reasoning methods that allow us to predict perturbation responses. We focus on gene regulatory networks (GRNs) as our knowledge bases, which contain activation (+) and inhibition (-) interaction relations between genes. We utilize symbolic reasoning via Boolean matrices (Ioannidis & Wong, 1991; Ai, 2025). Direct activation and inhibition interactions are compiled as $n\times n$ adjacency matrices $\langle{\bm{R}}_{+}^{(0)},{\bm{R}}_{-}^{(0)}\rangle$:
@@ 2.3 Abductive Learning
A general ABL training pipeline takes a neural model $f$ pretrained on labelled data $\langle{\bm{X}}_{l},{\bm{Y}}_{l}\rangle$ as initialization. From the unlabelled dataset ${\bm{X}}_{u}$, $f$ makes neural predictions $\hat{{\bm{y}}}=f({\bm{x}}_{u})$, which may be inconsistent with $\mathcal{KB}$. Consistency optimization is then performed, with revising $\hat{{\bm{y}}}$ to $\bar{{\bm{y}}}$ and updating $f$ on the revised dataset $\langle{\bm{X}}_{u},\bar{{\bm{Y}}}\rangle$. This process can be executed iteratively until convergence or reaching an iteration limit $T$. Formally,
$f:{\bm{X}}_{l}\rightarrow{\bm{Y}}_{l}$
$\mathrm{s.t.}\quad$ $\forall{\bm{x}}\in{\bm{X}}_{l}\cup{\bm{X}}_{u}:\ \mathcal{KB}\models\langle{\bm{x}},f({\bm{x}})\rangle,\ \mathrm{or}$
$f({\bm{x}})=\delta_{\mathcal{KB}}({\bm{x}}),\mathcal{KB}\models\langle{\bm{x}},\delta_{\mathcal{KB}}({\bm{x}})\rangle$
where $\delta({\bm{x}},\mathcal{KB})$ is the symbolic prediction on $\bm{x}$} by $\mathcal{KB}$ and $\models$” denotes logical entailment.
@@ 3.1 The Balanced Consistency Metric
To evaluate the consistency of a prediction against both the test data and the knowledge base, we define a balanced consistency metric $F_{1\ \mathrm{balance}}$, which considers the $F_{1}$ scores from both the test dataset and the knowledge base. $F_{1\ \mathrm{balance}}$ includes a coefficient $\gamma>1$ to balance the two $F_{1}$ scores and penalize when either score being too low:
@@ 3.2 ALIGNED Framework Overview
The ALIGNED framework (Figure 2) integrates three components to balance data-knowledge inconsistencies through iterative refinement. The neural component $f_{y}$ is a neural network which predicts perturbation responses from input data, while the symbolic component $\mathcal{KB}$ performs symbolic reasoning over gene regulatory networks encoded as matrices (computed via Equation 2.2). The adaptor $f_{a}$ learns to combine neural and symbolic predictions based on their relative reliability for each prediction.
@@ 3.3 Adaptive Neuro-Symbolic Alignment with Gradient-Free Optimization
In this section, we will introduce the alignment mechanism used by ALIGNED to adaptively integrate neural and symbolic predictions. We denote a binary alignment indicator vector ${\bm{a}}=f_{a}({\bm{x}})$ and the neuro-symbolic prediction $\bar{{\bm{y}}}$. After the initialization and each round of bidirectional update, $\bar{{\bm{y}}}$ is produced from both neural prediction $\hat{{\bm{y}}}$ and symbolic prediction $\delta_{\mathcal{KB}}({\bm{x}})$, according to the indicator $\bm{a}$} such that neural prediction is used when ${a}_{i}=0$ and symbolic prediction when ${a}_{i}=1$.
Our definition of the training objectives for the adaptor can be divided into three parts. First, since information derived from experimental data and curated knowledge may be inconsistent, the adaptor considers how neural-symbolic predictions differ from the curated knowledge. We describe this using the inconsistency between $\bar{{\bm{y}}}$ and $\mathcal{KB}$ based on Equation 2:
@@
Second, we design a loss term to leverage as much information from labelled training data as possible to reduce the predictions’ inconsistency with data. Therefore, we restrict the framework to only use knowledge-derived information when necessary. We defined this restriction with threshold $\theta$:
@@
Third, we take into account how well each gene is represented in knowledge. To measure this, we use a weight vector $\bm{w}$} as hyper-parameter, which contains the number of training data samples that are inconsistent with $\mathcal{KB}$ (computed by Equation 2), and the number of annotations from Gene Ontology (Ashburner et al., 2000). This allows us to reward $\bm{a}$} by maximizing the usage of symbolic prediction when a gene is represented well in $\mathcal{KB}$, otherwise use neural prediction. We define a loss with regard to $\bm{w}$}:
@@
where hyper-parameter $C_{w}$, $C_{l}$ are trade-off coefficients. Minimizing $L_{a}$ includes querying the symbolic $\mathcal{KB}$, which has a discrete structure. This creates a combinatorial optimization problem, so a gradient-free optimization method is necessary. We train $f_{a}$ with the REINFORCE algorithm (Williams, 1992; Hu et al., 2025) and initialize its sampling distribution based on $\bm{w}$} to reduce sampling complexity. To exploit representations captured by the neural component $f_{y}$ from the experimental data, $f_{a}$ shares input $\bm{x}$} and embedding layers with $f_{y}$. We optimize $f_{y}$ and $f_{a}$ jointly with the following objective:
@@ 3.4 Gradient-Based Knowledge Refinement with Sparse Regularization
To address missing and inaccurate interactions in $\mathcal{KB}$, we incorporate a knowledge refinement mechanism into the ALIGNED framework that leverages reliable information from neural and symbolic predictions. For computational efficiency on large-scale GRNs, we consider gradient-based optimization, and introduce an approximation function $\varepsilon(\cdot)$ for Boolean elements (Ravanbakhsh et al., 2016). This approximation enables gradient-based optimization compatibility of the non-differentiable Boolean matrix multiplication in Equation 2.2:
@@
We introduce an inductive bias for minimal modifications to the GRN during refinement. This ensures the biological relationships and structure in the GRN are not distorted by noise in the data. We perform an $l_{1}$ sparse regularizied optimization to achieve this, fitting $\mathcal{KB}$ to neuro-symbolic predictions using proximal gradient descent (Tibshirani, 1996; Candes & Recht, 2012). The objective of knowledge refinement is defined as follows:
$\min\limits_{{\bm{P}}^{(0)}_{+},{\bm{P}}^{(0)}_{-}\in\mathbb{R}_{+}^{n\times n}}\ \mathcal{L}_{\mathrm{refine}}({\bm{P}}^{(0)}_{+},{\bm{P}}^{(0)}_{-},k)=$ $\sum\limits_{{\bm{x}},{\bm{y}}\in\langle{\bm{X}}_{u},\bar{{\bm{Y}}}\rangle}\|\varepsilon_{t_{k}}({\bm{P}}_{+}^{(k)}-{\bm{P}}_{-}^{(k)})^{\top}{\bm{x}}-{\bm{y}}\|_{2}^{2}$
$+\lambda\big(\|\varepsilon_{t_{0}}({\bm{P}}_{+}^{(0)})-{\bm{R}}_{+}^{(0)}\|_{1}+\|\varepsilon_{t_{0}}({\bm{P}}_{-}^{(0)})-{\bm{R}}_{-}^{(0)}\|_{1}\big)$ (6)
where ${\bm{R}}^{(0)}_{+}$ and ${\bm{R}}^{(0)}_{-}$ represent the initial GRN before refinement, real-valued non-negative matrices ${\bm{P}}_{+}^{(0)}$ and ${\bm{P}}_{-}^{(0)}$ are the refined GRN with direct regulatory interactions. To facilitate gradient passing, we use real-number matrix computation instead of Boolean matrix in Equation 2.2 and use ${\bm{P}}_{+}^{(0)}$ and ${\bm{P}}_{-}^{(0)}$ to compute indirect regulatory interactions ${\bm{P}}_{+}^{(k)}$ and ${\bm{P}}_{-}^{(k)}$. $\lambda$ denotes a regularization parameter, $t_{k},t_{0}$ denote coefficients of approximation, $\|\cdot\|_{1}$ denotes the element-wise matrix $l_{1}$ norm.
@@ 4 Experiments
We evaluate ALIGNED on multiple large-scale perturbation datasets for predicting genome-wide responses and assess the knowledge refinement mechanism in isolation. The experiments address the following research questions:
@@ 4.1 Perturbation Prediction on Benchmark Datasets (Q1)
We focused on multiple large-scale perturbation datasets that are widely adopted for this prediction task: 1) Norman et al. (2019) for human K562 cells, including gene expression profiles under single and double perturbations across 102 genes (128 double and 102 single perturbations) with 89,357 samples; 2) Dixit et al. (2016) for mouse BDMC cells, containing 19 single gene perturbations with 43,401 samples; and 3) Adamson et al. (2016) for human K562 cells, containing 82 single gene perturbations with 65,899 samples. Our knowledge base $\mathcal{KB}$ integrates the Omnipath GRN (Türei et al., 2016) and the GO-based gene interaction graph from Roohani et al. (2024), covering 3,949 genes for the Norman et al. (2019) dataset and 2,958 genes for the Dixit et al. (2016) dataset. To evaluate methods on unseen perturbations, we split the test set of Norman et al. (2019) dataset to include 19 unseen single-gene perturbations and 18 unseen double-gene perturbations. The Dixit et al. (2016) and Adamson et al. (2016) datasets were split randomly.
We evaluated ALIGNED variants built with an MLP or a $\mathcal{KB}$-embedded GNN as the neural component. Performance of ALIGNED was compared in Figure 3 with state-of-the-art methods including: 1) GEARS, a GNN-based data-knowledge hybrid model (Roohani et al., 2024); 2) foundation models scGPT (Cui et al., 2024) and scFoundation Hao et al. (2024); 3) a linear additive perturbation model incorporating regulatory knowledge (Ahlmann-Eltze et al., 2025). To ensure that all methods are measured on the same knowledge base, ALIGNED does not perform knowledge refinement during this comparison.
To assess the contribution of each framework component, we conducted an ablation study comparing the complete ALIGNED framework against its neural component baseline (trained only on labelled data) in Figure 4. We tracked performance through two complete iterations of ALIGNED, with each iteration consisting of adaptive neuro-symbolic alignment followed by knowledge refinement (denoted as Align/Refine 1, 2).
For each method, we evaluated data consistency $F_{1}(\bar{{\bm{Y}}},{\bm{Y}}_{test})$ which measures performance of the predictions, knowledge consistency $F_{1}(\bar{{\bm{Y}}},\delta_{\mathcal{KB}}({\bm{X}}))$, and the balanced consistency metric $F_{1\ \mathrm{balance}}$ as defined in Equation 3.
Balanced Cons.Data Cons.Knowledge Cons.0.30.350.40.450.50.55(a) Norman et al. DatasetBalanced Cons.Data Cons.Knowledge Cons.0.10.20.30.40.5(b) Dixit et al. DatasetBalanced Cons.Data Cons.Knowledge Cons.0.20.30.40.50.6(c) Adamson et al. DatasetALIGNED (GNN)ALIGNED (MLP)LinearGEARSscGPTscFoundation
@@
[…]
Figure 4: Performance of the complete ALIGNED framework built with GNN.

### v2

@@ 1 Introduction
Two complementary approaches have emerged, either by leveraging latent representations trained on extensive cell data (Lotfollahi et al., 2023; Theodoris et al., 2023; Cui et al., 2024; Hao et al., 2024) or incorporating prior biological knowledge for inductive biases (Roohani et al., 2024; Wang et al., 2024; Littman et al., 2025; Wenkel et al., 2025). Yet, both approaches provide limited insights into the biological mechanisms underlying their predictions. Data-driven models operate as black boxes, making it difficult to understand which regulatory relationships drive specific predictions (Bendidi et al., 2024). While hybrid methods incorporate prior biological knowledge, they treat this knowledge as static constraints rather than interpretable and updatable representations of biological understanding. Importantly, current approaches provide no end-to-end solution to identify and resolve divergences between data-driven learning and existing knowledge, which limits opportunities for continual refinement of biological understanding (Gavriilidis et al., 2024; Zuidberg Dos Martires et al., 2024; Kedzierska et al., 2025).
@@
  • Adaptive Neuro-Symbolic Alignment. We adaptively combine neural and symbolic predictions from inconsistent information sources by weighting neural and symbolic components with a gradient-free optimization mechanism.
@@
  • Results. ALIGNED matches or exceeds state-of-the-art methods in predictive accuracy while substantially outperforms existing methods in balanced consistency. In addition, we show that ALIGNED’s knowledge refinement can re-discover cross-referenced regulatory relationships.
@@ 2.1 Problem Setting
We formalize the prediction of genome-scale response to genetic perturbation as a ternary classification problem. The goal is to learn a function $f:\{-1,0,1\}^{n}\rightarrow\{-1,0,1\}^{n}$, where $n$ is the total number of genes. The input values 1-1, 0, and $1$ represent negative perturbation (deletion or knockout), no perturbation, and positive perturbation (overexpression), respectively. The output values indicate decreased expression, no significant change, or increased expression for each gene. Ternary representation of gene expression profiles is commonly used in downstream biological research tasks, such as gene differential-expression analysis and pathway annotations (Love et al., 2014; Badia-i Mompel et al., 2023).
@@ 2.2 Symbolic reasoning over gene regulatory networks
We focus on gene regulatory networks (GRNs) as our knowledge bases, which contain activation (+) and inhibition (-) interaction relations between genes. It is a widely used qualitative approach to capture directed regulatory effects (Chevalley et al., 2025). We utilize symbolic reasoning via Boolean matrices (Ioannidis & Wong, 1991; Ai, 2025). Direct activation and inhibition interactions are compiled as $n\times n$ adjacency matrices $\langle{\bm{R}}_{+}^{(0)},{\bm{R}}_{-}^{(0)}\rangle$:
@@ 2.3 Abductive Learning
A general ABL training pipeline takes a neural model $f$ pretrained on labelled data $\langle{\bm{X}}_{l},{\bm{Y}}_{l}\rangle$ as initialization. From the unlabelled dataset ${\bm{X}}_{u}$, $f$ makes neural predictions $\hat{{\bm{y}}}=f({\bm{x}}_{u})$, which may be inconsistent with $\mathcal{KB}$. Consistency optimization is then performed, with revising $\hat{{\bm{y}}}$ to $\bar{{\bm{y}}}$ and updating $f$ on the revised dataset $\langle{\bm{X}}_{u},\bar{{\bm{Y}}}\rangle$. This process can be executed iteratively until convergence or reaching an iteration limit $T$. Formally, the objective of ABL is to make predictions that are most consistent with $\mathcal{KB}$:
$\max\limits_{f}\ p({\bm{y}}\mid{\bm{x}};f,\mathcal{KB})=\mathbb{I}\big(\mathcal{KB}\cup{\bm{x}}\models f({\bm{x}})\big)\ p({\bm{y}}\mid{\bm{x}};f)$
where on the right hand side $\models$” denotes logical entailment, the first term is an indicator of consistency and the second term denotes the posterior distribution of neural predictions.
@@ 3.1 The Balanced Consistency Metric
To evaluate the consistency of a prediction against both the test data and the knowledge base, we define a balanced consistency metric $F_{1\ \mathrm{balance}}$, which considers the macro $F_{1}$ scores from both the test dataset and the knowledge base. $F_{1\ \mathrm{balance}}$ includes a coefficient $\gamma>1$ to balance the two macro $F_{1}$ scores and penalize when either score being too low:
@@
This metric evaluates the model’s performance on making generalizable predictions while remaining grounded in curated biological prior. Unless specified, all $F_{1}$ scores in evaluations are computed as macro $F_{1}$.
@@ 3.2 ALIGNED Framework Overview
The ALIGNED framework (Figure 2) integrates three components to leverage reliable information from data and knowledge for predicting perturbation response and refining regulatory knowledge. The neural component $f_{y}$ is a neural network which predicts perturbation responses from input data, while the symbolic component $\mathcal{KB}$ performs symbolic reasoning over gene regulatory networks encoded as matrices (computed via Equation 2.2). The adaptor $f_{a}$ learns to combine neural and symbolic predictions based on their relative reliability for each prediction. ALIGNED assumes that (i) when perturbation data is relatively more reliable than prior knowledge, the adaptor would increase reliance on the neural component, (ii) when the prior knowledge provides more reliable information, the symbolic component would contribute more to overall predictions.
@@ 3.3 Adaptive Neuro-Symbolic Alignment with Gradient-Free Optimization
In this section, we will introduce the alignment mechanism used by ALIGNED to adaptively integrate neural and symbolic predictions. We denote a binary alignment indicator vector ${\bm{a}}=f_{a}({\bm{x}})$ and the neuro-symbolic prediction $\bar{{\bm{y}}}$. After the initialization and each round of bidirectional update, $\bar{{\bm{y}}}$ is produced from both neural prediction $\hat{{\bm{y}}}$ and symbolic prediction $\delta_{\mathcal{KB}}({\bm{x}})$ according to the indicator $\bm{a}$}, such that $\bar{{y}}_{i}=\begin{cases}\hat{{y}}_{i},\ &{a}_{i}=0\\ \delta_{\mathcal{KB}}({\bm{x}})_{i},\ &{a}_{i}=1\end{cases}$.
Our definition of the training objectives for the adaptor can be divided into three parts. First, since information derived from experimental data and curated knowledge may be inconsistent, the adaptor considers how neuro-symbolic predictions differ from the curated knowledge. We describe this using the inconsistency between $\bar{{\bm{y}}}$ and $\mathcal{KB}$ based on Equation 2:
@@
Second, we design a loss term to leverage as much information from labelled training data as possible to reduce the predictions’ inconsistency with data. Therefore, we restrict the framework to only use knowledge-derived information when necessary. We defined the knowledge usage restriction with threshold $\theta$:
@@
Third, we take into account how well each gene is represented in knowledge. To measure this, we use a weight vector $\bm{w}$} as hyperparameter, which contains the number of training data samples that are inconsistent with $\mathcal{KB}$ (computed by Equation 2), and the number of annotations from Gene Ontology (Ashburner et al., 2000). This allows us to reward $\bm{a}$} by maximizing the usage of symbolic prediction when a gene is represented well in $\mathcal{KB}$, otherwise use neural prediction. We define a loss with regard to $\bm{w}$}:
@@
where hyperparameter $C_{w}$, $C_{l}$ are trade-off coefficients. Minimizing $L_{a}$ includes querying the symbolic $\mathcal{KB}$, which has a discrete structure. This creates a combinatorial optimization problem, so a gradient-free optimization method is necessary. We train $f_{a}$ with the REINFORCE algorithm (Williams, 1992; Hu et al., 2025) and initialize its sampling distribution based on $\bm{w}$} to reduce sampling complexity. To exploit representations captured by the neural component $f_{y}$ from the experimental data, $f_{a}$ shares input $\bm{x}$} and embedding layers with $f_{y}$. Adaptive alignment optimizes $f_{y}$ and $f_{a}$ jointly with the following objective:
@@ 3.4 Gradient-Based Knowledge Refinement with Sparse Regularization
To address missing and inaccurate interactions in $\mathcal{KB}$, we incorporate a knowledge refinement mechanism into the ALIGNED framework that leverages reliable information from neural and symbolic predictions. For computational efficiency on large-scale GRNs, we consider gradient-based optimization, and introduce an approximation function $\varepsilon(\cdot)$ for Boolean elements (Ravanbakhsh et al., 2016). This approximation enables gradient-based optimization compatibility of the non-differentiable Boolean matrix operations in Equation 2.2:
@@
We introduce an inductive bias for minimal modifications to the GRN during refinement. This ensures the biological relationships and structure in the GRN are not distorted by spurious signals from data, such as overfitting to noise or learning shortcut patterns. We perform an $l_{1}$ sparse regularized optimization to achieve this, fitting $\mathcal{KB}$ to neuro-symbolic predictions using proximal gradient descent (Tibshirani, 1996; Candes & Recht, 2012). Knowledge refinement optimizes the following objective:
$\min\limits_{\bar{{\bm{R}}}^{(0)}_{+},\bar{{\bm{R}}}^{(0)}_{-}\in\mathbb{R}_{+}^{n\times n}}\ \mathcal{L}_{\mathrm{refine}}(\bar{{\bm{R}}}^{(0)}_{+},\bar{{\bm{R}}}^{(0)}_{-},k)=$ $\sum\limits_{{\bm{x}},{\bm{y}}\in\langle{\bm{X}}_{u},\bar{{\bm{Y}}}\rangle}\|\varepsilon_{t_{k}}(\bar{{\bm{R}}}_{+}^{(k)}-\bar{{\bm{R}}}_{-}^{(k)})^{\top}{\bm{x}}-{\bm{y}}\|_{2}^{2}$
$+\lambda\big(\|\varepsilon_{t_{0}}(\bar{{\bm{R}}}_{+}^{(0)})-{\bm{R}}_{+}^{(0)}\|_{1}+\|\varepsilon_{t_{0}}(\bar{{\bm{R}}}_{-}^{(0)})-{\bm{R}}_{-}^{(0)}\|_{1}\big)$ (6)
where $\mathcal{KB}$ is substituted with $\langle\varepsilon_{t_{k}}(\bar{{\bm{R}}}_{+}^{(k)}),\varepsilon_{t_{k}}(\bar{{\bm{R}}}_{-}^{(k)})\rangle$ after the knowledge refinement step.
${\bm{R}}^{(0)}_{+}$ and ${\bm{R}}^{(0)}_{-}$ represent the initial GRN before refinement, real-valued non-negative matrices $\bar{{\bm{R}}}_{+}^{(0)}$ and $\bar{{\bm{R}}}_{-}^{(0)}$ are the refined GRN with direct regulatory interactions. To facilitate gradient passing, we use real-number matrix computation instead of Boolean matrix in Equation 2.2 and use $\bar{{\bm{R}}}_{+}^{(0)}$ and $\bar{{\bm{R}}}_{-}^{(0)}$ to compute indirect regulatory interactions $\bar{{\bm{R}}}_{+}^{(k)}$ and $\bar{{\bm{R}}}_{-}^{(k)}$. $\lambda$ denotes a regularization parameter, $t_{k},t_{0}$ denote coefficients of approximation, $\|\cdot\|_{1}$ denotes the element-wise matrix $l_{1}$ norm.
@@ 4 Experiments
We evaluate ALIGNED on multiple large-scale perturbation datasets for predicting genome-wide responses and assess the knowledge refinement mechanism in isolation [^1] Details of experiment hardware and runtime are in Section E.2.. The experiments address the following research questions:
@@ 4.1 Perturbation Prediction on Benchmark Datasets (Q1)
We focused on multiple large-scale perturbation datasets that are widely adopted for this prediction task: (i) Norman et al. (2019) for human K562 cells, including gene expression profiles under single and double perturbations across 102 genes (128 double and 102 single perturbations) with 89,357 samples; (ii) Dixit et al. (2016) for mouse BDMC cells, containing 19 single gene perturbations with 43,401 samples; and (iii) Adamson et al. (2016) for human K562 cells, containing 82 single gene perturbations with 65,899 samples. ALIGNED predicts perturbation response profiles for all genes in the datasets. Our knowledge base $\mathcal{KB}$ integrates the Omnipath GRN (Türei et al., 2016) and the GO-based gene interaction graph from Roohani et al. (2024), covering 3,949 genes for the Norman et al. (2019) dataset and 2,958 genes for the Dixit et al. (2016) dataset. To evaluate methods on unseen perturbations, we split the test set of Norman et al. (2019) dataset to include 19 unseen single-gene perturbations and 18 unseen double-gene perturbations. The Dixit et al. (2016) and Adamson et al. (2016) datasets were split randomly.
We evaluated ALIGNED variants built with an MLP or a $\mathcal{KB}$-embedded GNN as the neural component. Performance of ALIGNED was compared in Figure 3 with state-of-the-art methods including: (i) GEARS, a GNN-based data-knowledge hybrid model (Roohani et al., 2024); (ii) foundation models scGPT (Cui et al., 2024) and scFoundation (Hao et al., 2024); (iii) a transformer-based model State (Adduri et al., 2025); (iv) a linear additive perturbation model incorporating regulatory knowledge (Ahlmann-Eltze et al., 2025). To ensure that all methods are measured on the same knowledge base, ALIGNED does not perform knowledge refinement during this comparison.
The ALIGNED framework with GNN and MLP results in Figure 3 performed adaptive neuro-symbolic alignment, but not knowledge refinement. For each method, we evaluated (i) data consistency $F_{1}(\bar{{\bm{Y}}},{\bm{Y}}_{test})$ which measures predictive performance; (ii) knowledge consistency $F_{1}(\bar{{\bm{Y}}},\delta_{\mathcal{KB}}({\bm{X}}))$ which quantifies the extent to which the model remains grounded in curated biological prior; and (iii) the balanced consistency metric $F_{1\ \mathrm{balance}}$ defined as Equation 3 evaluating the trade-off between inconsistent sources.
Balanced Cons.Data Cons.Knowledge Cons.0.20.30.40.5(a) Norman et al. DatasetBalanced Cons.Data Cons.Knowledge Cons.0.10.20.30.40.5(b) Dixit et al. DatasetBalanced Cons.Data Cons.Knowledge Cons.0.20.30.40.50.6(c) Adamson et al. DatasetALIGNED (GNN)ALIGNED (MLP)LinearGEARSStatescGPTscFoundation